Summary data on daily caffeine consumption for a sample of adult women is the following: n = 36, sample mean= 215 mg. Assume the population is normally distributed with standard deviation is σ = 12 mg.

Construct an appropriate 98% confidence interval for the population mean, μ.

A.) What is your point estimate?

B.) Critical Value?

C.) Standard Error?

D.) Margin of Error?

E.) Lower Confidence Bound (or limit)?

F.) Upper Confidence Bound (or limit)?
Answer #1

Given, n-36 (1-0)100 =98 X = 215 I-d = 98/100 J12 1-K= 0.98 do 1-0.98 -0.02 0.02 -0.01 0.02 2 (A) sample mean = 215 is pointof error is 4.652 The Margin (E) lower confidence limit is lower confidence limit = x - E = 215 - 4.65 2 = 210.348 is 210.348